media%2F06f%2F06fe0dce-6894-4dd2-bf95-e2
Here again is the histogram showing the distribution of times, in minutes, required for 25 rats in an animal behavior experiment to navigate a maze successfully Time for a Group of Rates to Navigate a Maze Time in Minutes Which of the following best describes the shape of the histogram? Symmetric Fught-skewed with no outliers Left-skewed with no outliers Right-skewed with a possible outlier Left-skewed with a possible outier
Here again is the histogram phowing the distribution of times, in minutes, required for 25 rats in an animal behavior experiment to navigate a maze successfully Time for a Group of Rates to Navigate a Maze Time in Minutes For the data described by the above histogram, (A) the median will be smaller than the mean (B) the median will be larger than the mean (C) the median and the mean will be about the same (D) the median and the range will be about the same. both (A) and (D) are corect ords
Answer #1

a) The percentage of the rate navigated the maze in less than 5.5 minutes is

[ (3+8+6+4) / (3+8+6+4+2+1+1)] * 100 = 0.84 * 100 = 84%

Correct answer: option (E) 84%

b) Correct ansewr: Option (D) Right-Skewed with a possible outlier

c) Correct answer: Option (A) the median will be smaller than the mean